A cone made of insulating material has total charge Q spread uniformly over its sloping surface. Calculate the work done in bringing a small test charge q from infinity to the apex of the cane. The cone has a slope length L.

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear ayush,



y/r = x/L => y= xr/L

dq = 2Πydxσ = 2Πrσxdx/l

for this ring potential at vertex = dV = ∫kdq/x

integrating from 0 - L ,
     
         V = 2Πkrσ = Q / 2Πε0L

Now work done = Vq = Qq / 2Πε0L

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delh